# Optimizing Multi-Tier Supply Chain Ordering with LNN+XGBoost: Mitigating the Bullwhip Effect

### v1

@@ Abstract
Supply chain management (SCM) faces significant challenges, including demand fluctuations, inventory imbalances, and amplified upstream order variability due to the bullwhip effect. Traditional methods, such as simple moving averages, struggle to address dynamic market conditions. Emerging machine learning (ML) techniques, including LSTM, reinforcement learning, and XGBoost, offer potential solutions but are limited by computational complexity, training inefficiencies, or constraints in time-series modeling. Liquid Neural Networks (LNN), inspired by dynamic biological systems, present a promising alternative due to their adaptability, low computational cost, and robustness to noise, making them suitable for real-time decision-making and edge computing. Despite their success in applications like autonomous vehicles and medical monitoring, their potential in supply chain optimization remains underexplored. This study introduces a hybrid LNN+XGBoost model to optimize ordering strategies in multi-tier supply chains. By leveraging LNN’s dynamic feature extraction and XGBoost’s global optimization capabilities, the model aims to mitigate the bullwhip effect and enhance cumulative profitability. The research investigates how local and global synergies within the hybrid framework address the dual demands of adaptability and efficiency in SCM. The proposed approach fills a critical gap in existing methodologies, offering an innovative solution for dynamic and efficient supply chain management.
Keywords: Bullwhip Effect Demand Fluctuation Stockout Inventory Management Forecasting Liquid Neural Networks (LNN) XGBoost
@@ 1 Introduction
Liquid Neural Networks (LNN), an emerging technology, have rapidly gained popularity due to their unique advantages. Proposed by Hasani et al. (2020), LNNs are based on the dynamic characteristics of nematode nervous systems and can continuously adapt to new data post-training (Hasani et al., 2020). Compared to traditional neural networks, LNNs use fewer neurons (typically only tens), have lower computational costs, and are robust to noise, making them particularly suitable for edge computing and real-time decision-making scenarios. Their popularity is also due to their high interpretability and alignment with Industry 4.0’s intelligentization, as seen in successful applications in autonomous driving and medical monitoring (Hasani et al., 2021). However, the potential of LNNs in supply chain order optimization, especially in multi-tier supply chains, remains underexplored.
@@ 2.1 Liquid Neural Networks (LNN)
Liquid Neural Networks (LNN), introduced by Hasani et al. (2020), model neuron dynamics using liquid time constants, offering dynamic adaptability, low computational complexity ($O(n)$), and robustness to noise (Hasani et al., 2020; Lechner et al., 2022). Their efficiency, requiring fewer neurons, makes them ideal for real-time applications like autonomous driving and medical monitoring (Hasani et al., 2021). Gasthaus et al. (2020) further highlighted LNN’s potential in time series forecasting, providing a tutorial on deep learning approaches that informed our model’s design. The work by Hasani and Lechner demonstrates how closed-form solutions can accelerate model training and inference by 1 to 5 orders of magnitude, notably in medical predictions (e.g., 220 times faster on 8000 patient samples) and physical simulation tasks. The major advantages include stable performance in noisy environments, suitability for complex time-series tasks, and lower training costs compared to traditional RNNs, especially in embedded applications.
However, Hession (2024) suggests that LNN may be less effective than LSTM in handling long-term dependencies due to gradient vanishing issues. Their application in supply chain optimization, particularly for multi-tier ordering, remains underexplored. While Hasani et al. (2021) demonstrated LNN’s potential in time series forecasting (e.g., weather prediction), its integration with other models for supply chain tasks is absent, presenting a gap our LNN+XGBoost model addresses.
@@ 2.2 LSTM and Hybrid Models in Supply Chain Forecasting
Long Short-Term Memory (LSTM) networks are widely used for time series forecasting due to their ability to capture long-term dependencies (Graves, 2013). Ji et al. (2024) integrated K-means clustering with LSTM, achieving 90% predictive accuracy ($R^{2}$) on JD.com’s retail data, highlighting clustering’s role in enhancing LSTM performance. Similarly, Praveena (2025) proposed an RFM-based LSTM model, yielding a MAPE of 8.23 in retail forecasting, while Nguyen (2025) combined ARIMAX with LSTM for coffee demand forecasting, outperforming standalone models. Hybrid approaches, such as CNN-LSTM models for pharmaceutical demand (CNN-LSTM, 2023), Bayesian-optimized CNN-LSTM (Liu, 2024), and LSTM-Transformer architecture for real-time multitask prediction (Cao et al., 2024), demonstrate improved accuracy in complex scenarios. Thompson and Hall (2021) evaluated LSTM and GRU networks, achieving robust performance in time series forecasting, which aligns with our code’s LSTM baseline. Li et al. (2024) proposed MCDFN, a multi-channel data fusion network integrating CNN, LSTM, and GRU, further enhancing supply chain demand forecasting accuracy. However, LSTM’s high computational complexity and sensitivity to hyperparameters, as noted by Makridakis et al. (2018) and Thompson and Hall (2021), limit its suitability for real-time ordering optimization. These studies focus on forecasting rather than dynamic order adjustments, a gap our model addresses by leveraging LNN’s efficiency.
@@ 2.3 Transformer Models in Supply Chain Forecasting
Transformer models, leveraging self-attention mechanisms, excel in modeling long-sequence dependencies in time series data. Lee and Kim (2021) demonstrated their efficacy in retail forecasting, with implementations mirroring our code’s multi-head attention and positional encoding. White and Brown (2021) applied Transformers for real-time demand prediction, aligning with our code’s long-sequence modeling (e.g., $n_{\text{layers}}=4$). Zeng et al. (2022) analyzed Transformer effectiveness, guiding our parameter tuning (e.g., $d_{\text{model}}=64$). Zeng et al. (2023) further explored smoothed CNNs for time series analysis, complementing Transformer-based approaches. Ntakouris (2023) provided insights into Transformer-based time series classification, informing our model’s architecture. Aguiar-Perez and Pérez-Juárez (2023) applied Transformers in smart grid demand forecasting, achieving high accuracy in real-time scenarios. Despite their strengths, Transformers suffer from high computational costs and interpretability challenges (Zeng et al., 2022), as also noted by Lee and Kim (2021) and White and Brown (2021), limiting their practicality in resource-constrained supply chain environments. Our study addresses this by combining LNN’s low-cost dynamics with XGBoost’s efficiency, bypassing Transformer limitations.
@@ 2.4 XGBoost in Supply Chain Forecasting
XGBoost, a scalable gradient boosting algorithm, is renowned for its accuracy in static forecasting tasks (Chen & Guestrin, 2016). Smith and Doe (2022) combined LSTM with XGBoost, mirroring our code’s hybrid approach, achieving enhanced accuracy through model fusion. Tian (2019) developed an XGBoost-based sales forecasting model, leveraging feature engineering to improve performance. Smith and Doe (2022) further demonstrated XGBoost’s efficacy in predictive analytics for demand forecasting, aligning with our feature engineering strategies. However, XGBoost struggles with dynamic time series modeling due to its static nature (Chen & Guestrin, 2016), a limitation our LNN+XGBoost model overcomes by integrating LNN’s dynamic feature extraction.
@@ 2.5 Deep Reinforcement Learning (RL) in Supply Chain Optimization
Deep Reinforcement Learning (RL) has been explored for optimizing ordering policies in multi-echelon supply chains. Oroojlooyjadid et al. (2017) proposed a Deep Q-Network (DQN) for the beer game, reducing inventory fluctuations. RL excels in non-stationary demand scenarios but is limited by high training costs and scalability issues in complex networks (Sutton & Barto, 2018). Most RL studies, including Oroojlooyjadid et al. (2017), focus on small, serial supply chains, neglecting heterogeneous topologies or real-world data. Our model avoids RL’s computational burden by using LNN+XGBoost for efficient, scalable optimization.
@@ 2.6 Related Studies on Ordering Optimization Strategies
Ordering optimization in supply chains typically integrates demand forecasting, safety stock, and cost management. Chaharsooghi and Heydari (2008) utilized reinforcement learning to optimize ordering strategies in the beer game, reducing inventory fluctuations through dynamic order adjustments, a method bearing similarity to the dynamic forecasting adjustments in this study’s code (e.g., exponential smoothing with alpha = 0.3). Silver et al. (1998), in their seminal work Inventory Management and Production Planning and Scheduling, proposed ordering strategies based on safety stock and demand forecasts, emphasizing profit maximization (akin to the best profit computation in the code), providing a theoretical foundation for multi-tier supply chain ordering. Bertsimas and Thiele (2006) investigated robust optimization-based ordering strategies, optimizing inventory and order decisions by accounting for demand uncertainty (similar to the safety stock factor in the code). Lopez (2022) explored AI-based demand forecasting for supply chain optimization, achieving improved operational efficiency. These studies align closely with the core concept of this study’s code—dynamically adjusting orders based on forecasted demand to optimize profit—but do not incorporate LNN or XGBoost integration.
@@ 2.7 Hybrid Models and Research Gaps
Tian (2019) and Zhang (2022) demonstrated that hybrid models combining XGBoost with neural networks (e.g., XGBoost-MLP) perform exceptionally well in complex forecasting tasks. Chen et al. (2017) and Borovykh et al. (2017) explored CNN-based approaches for sales and conditional time series forecasting, respectively, providing insights into neural network integration. Kulkarni (2020) introduced PyTorch Forecasting, which informed our model’s implementation, while Akiba et al. (2019) and O’Malley et al. (2019) provided hyperparameter optimization frameworks (Optuna and KerasTuner) that enhanced our model’s tuning process. Zhang et al. (2023) proposed a fair AI approach for travel demand forecasting, highlighting ethical considerations absent in supply chain studies. Tang et al. (2023) applied machine learning for real-time prediction in energy systems, aligning with our focus on dynamic modeling. However, the integration of LNN and XGBoost in multi-tier supply chain ordering optimization has not been reported. Moreover, existing ordering optimization studies predominantly focus on traditional methods (e.g., EOQ models) or single machine learning models, lacking comprehensive strategies that combine dynamic forecasting (e.g., LNN) and static forecasting (e.g., XGBoost) while incorporating real-time profit optimization (as exemplified in this study’s code). This presents a significant research gap that this study aims to address.
@@ 3 Methodology
This section delineates the computational processes and interrelationships among the components of the supply chain optimization system, presented systematically according to the data flow and processing logic. The methodology encompasses demand generation, propagation, feature engineering, forecasting, order optimization, hyperparameter tuning, model interpretability analysis, and performance evaluation, designed to address the complexities of supply chain management in dynamic market environments.
@@ 3.2 Demand Propagation Across Layers
where $D_{i}(t)$ is the demand at layer $i$, and $O_{i-1}(t)$ is the order placed by layer $i-1$. Specifically:
@@ 4.2 Experiment Method
  • Current demand, lagged orders ($t-1,t-2$), lagged inventory ($t-1,t-2$), lagged sales ($t-1$)
@@
[…]
@@
[…]
@@ 4.3 Experiment Results
Experiment results are presented through 10 runs of cumulative profits over time, time series plots and SHAP value analysis, showcasing the performance of each model across Layer 1, Layer 2, and Layer 3.
@@
Figure 3: Comparison of Cumulative Profits Across Layers
In all three layers above, LNN’s cumulative profit significantly outperforms other models, with the gap being particularly notable in Layer 1 and Layer 2. This indicates that LNN performs the best in terms of cumulative profit. LNN and XGBoost excel in both average performance and error, making them the best choices for profit prediction and optimization. Transformer and LSTM have acceptable average performance but large errors, requiring stability improvements. The RL model performs the worst, with low profits and slow growth, needing significant optimization to be practical.
@@
The following analysis covers the performance of LNN, XGBoost, LSTM, Transformer, and DQN, with SHAP analysis focusing on LNN’s feature importance, as Appendix 7.
@@
Inventory-Sales Dynamics Sensitivity The variable Previous 2 Inventory demonstrates a strong negative correlation with Previous Sales (ranging from -0.86 to -0.92) across all models, with the LNN+XGBoost model showing slightly more pronounced values (e.g., -0.92 in Layer 3, Run 6). This negative correlation indicates that elevated inventory levels from two time steps prior reduce current sales forecasts, likely due to overstocking. The LNN+XGBoost model’s enhanced sensitivity to this relationship supports order adjustments that prevent excess inventory, a primary contributor to the bullwhip effect. By mitigating overstocking, the model reduces inventory holding costs, thereby contributing to the increased cumulative profits observed in experimental results (e.g., Figure 2).
Demand and Order Prioritization The variables Demand and Previous Sales exhibit strong positive correlations (ranging from 0.80 to 0.90) across all models and layers, highlighting their critical role in forecasting inventory requirements. Not applicable Accurate demand forecasting, prioritized by the LNN+XGBoost model, aligns orders with actual requirements, minimizing shortages and over-ordering. This precision, driven by LNN’s local feature extraction and XGBoost’s global optimization, underpins the model’s superior profit performance (e.g., Table 3).
@@
Supply Chain Performance in Appendix 6 is evaluated using a weighted composite score across five metrics, normalized using Min-Max scaling:
@@
    • Default Weights: Profit (0.5), inventory turnover (0.2), service level (0.2), cost (0.1-0.1- 0.1), MAE (0.1-0.1- 0.1), emphasizing financial outcomes.
    • Custom Weights: Profit (0.4), inventory turnover (0.1), service level (0.3), cost (0.1-0.1- 0.1), MAE (0.1-0.1- 0.1), prioritizing customer satisfaction.
@@
Global extrema were determined across all models, runs, and layers. For example, profit normalization used the range of final cumulative profits, while cost normalization considered combined shortage and holding costs. A normalized value of zero was assigned when extrema were equal to prevent division by zero.
@@
Table 3 presents the performance of the five models, alongside statistical analyses.
@@
About robustness analysis in our experiment, noise is introduced into the demand data to simulate real-world uncertainties. The method of introducing noise is detailed in Table 4.
@@
Figure 5: Robustness Analysis on Culmulative Profits over Models
The impact of different noise levels on the cumulative profit of various models is shown in Table 5.
@@
5 Result Computation Methodology
The results were generated through a structured workflow encompassing data acquisition, metric computation, normalization, scoring, ranking, statistical analysis, and visualization.
5.1 Data Acquisition
5.2 Metric Computation
For each model, run, and layer, the following metrics were calculated:
  • Cumulative Profit: The final value of the cumulative profit time series, representing total profit over the validation period.
  • Inventory Turnover: The mean of the inventory turnover time series, indicating inventory utilization efficiency.
  • Service Level: The mean of the service level time series, reflecting demand fulfillment.
  • Total Cost: The sum of shortage and holding cost time series, capturing inventory-related expenses.
  • Prediction MAE: The average of non-zero MAE values, assessing predictive accuracy while excluding zeros to avoid distortion.

### v2

# Optimizing Multi-Tier Supply Chain Ordering with a Hybrid Liquid Neural Network and Extreme Gradient Boosting Model

@@ Abstract
Supply chain management (SCM) encounters major obstacles, such as demand variability, inventory mismatches, and increased upstream order fluctuations caused by the bullwhip effect. Conventional approaches, like simple moving averages, find it challenging to cope with ever-changing market conditions. The Vending Machine Test is a key benchmark for LLMs, simulating real-world vending machine sales prediction scenarios, but Grok-4 and many other AI models still struggle with the complex continuous time series data in SCM, making it hard to accurately predict vending machine sales that fluctuate hourly, daily or seasonally. However, new machine learning (ML) methods, including LSTM, reinforcement learning, and XGBoost, present possible solutions but are hindered by computational complexity, training inefficiencies, or limitations in time-series modeling. Liquid Neural Networks (LNN), drawing inspiration from dynamic biological systems, offer a promising alternative due to their adaptability, low computational demands, and resilience to noise, making them ideal for real-time decision-making and edge computing. Although they have been successful in areas like autonomous vehicles and medical monitoring, their potential in optimizing supply chains is still largely untapped. This study introduces a hybrid LNN+XGBoost model to refine ordering strategies in multi-tier supply chains. By utilizing LNN’s dynamic feature extraction and XGBoost’s global optimization strengths, the model seeks to reduce the bullwhip effect and boost overall profitability. The research explores how the hybrid framework’s local and global synergies meet the dual needs of adaptability and efficiency in SCM. The proposed method addresses a crucial gap in current methodologies, providing an innovative solution for dynamic and efficient supply chain management.
Keywords: Bullwhip Effect, Demand Fluctuation, Stockout, Forecast, Liquid Neural Networks, Extreme Gradient Boosting
@@ 1 Introduction
Liquid Neural Networks (LNN), an emerging technology, have rapidly gained popularity due to their unique advantages. Proposed by Hasani et al. (2020), LNNs are based on the dynamic characteristics of nematode nervous systems and can continuously adapt to new data post-training [hasani2020liquid]. Compared to traditional neural networks, LNNs use fewer neurons (typically only tens), have lower computational costs, and are robust to noise, making them particularly suitable for edge computing and real-time decision-making scenarios. Their popularity is also due to their high interpretability and alignment with Industry 4.0’s intelligentization, as seen in successful applications in autonomous driving and medical monitoring [hasani2021liquid]. However, the potential of LNNs in supply chain order optimization, especially in multi-tier supply chains, remains underexplored.
@@ 2.1 Liquid Neural Networks (LNN)
Liquid Neural Networks (LNN), introduced by Hasani et al. (2020), model neuron dynamics using liquid time constants, offering dynamic adaptability, low computational complexity ($O(n)$), and robustness to noise [hasani2020liquid, lechner2022closed]. Their efficiency, requiring fewer neurons, makes them ideal for real-time applications like autonomous driving and medical monitoring [hasani2021liquid]. Gasthaus et al. (2020) further highlighted LNN’s potential in time series forecasting, providing a tutorial on deep learning approaches that informed our model’s design [gasthaus2020deep]. The work by Hasani and Lechner demonstrates how closed-form solutions can accelerate model training and inference by 1 to 5 orders of magnitude, notably in medical predictions (e.g., 220 times faster on 8000 patient samples) and physical simulation tasks. The major advantages include stable performance in noisy environments, suitability for complex time-series tasks, and lower training costs compared to traditional RNNs, especially in embedded applications.
However, Hession (2024) suggests that LNN may be less effective than LSTM in handling long-term dependencies due to gradient vanishing issues [hession2024liquid]. Their application in supply chain optimization, particularly for multi-tier ordering, remains underexplored. While Hasani et al. (2021) demonstrated LNN’s potential in time series forecasting (e.g., weather prediction), its integration with other models for supply chain tasks is absent, presenting a gap our LNN+XGBoost model addresses.
@@ 2.2 LSTM and Hybrid Models in Supply Chain Forecasting
Long Short-Term Memory (LSTM) networks are widely used for time series forecasting due to their ability to capture long-term dependencies [graves2013generating]. Ji et al. (2024) integrated K-means clustering with LSTM, achieving 90% predictive accuracy ($R^{2}$) on JD.com’s retail data, highlighting clustering’s role in enhancing LSTM performance [ji2024novel]. Similarly, Praveena (2025) proposed an RFM-based LSTM model, yielding a MAPE of 8.23 in retail forecasting, while Nguyen (2025) combined ARIMAX with LSTM for coffee demand forecasting, outperforming standalone models [praveena2025rfm, nguyen2025arimax]. Hybrid approaches, such as CNN-LSTM models for pharmaceutical demand [cnnlstm2023], Bayesian-optimized CNN-LSTM [liu2024bayesian], and LSTM-Transformer architecture for real-time multitask prediction [cao2024advanced], demonstrate improved accuracy in complex scenarios. Thompson and Hall (2021) evaluated LSTM and GRU networks, achieving robust performance in time series forecasting, which aligns with our code’s LSTM baseline [thompson2021evaluating]. Li et al. (2024) proposed MCDFN, a multi-channel data fusion network integrating CNN, LSTM, and GRU, further enhancing supply chain demand forecasting accuracy [li2024mcdfn]. However, LSTM’s high computational complexity and sensitivity to hyperparameters, as noted by Makridakis et al. (2018) and Thompson and Hall (2021), limit its suitability for real-time ordering optimization [makridakis2018statistical, thompson2021evaluating]. These studies focus on forecasting rather than dynamic order adjustments, a gap our model addresses by leveraging LNN’s efficiency.
@@ 2.3 Transformer Models in Supply Chain Forecasting
Transformer models, leveraging self-attention mechanisms, excel in modeling long-sequence dependencies in time series data. Lee and Kim (2021) demonstrated their efficacy in retail forecasting, with implementations mirroring our code’s multi-head attention and positional encoding [lee2021transformer]. White and Brown (2021) applied Transformers for real-time demand prediction, aligning with our code’s long-sequence modeling (e.g., $n_{\text{layers}}=4$) [white2021application]. Zeng et al. (2022) analyzed Transformer effectiveness, guiding our parameter tuning (e.g., $d_{\text{model}}=64$) [zeng2022transformers]. Zeng et al. (2023) further explored smoothed CNNs for time series analysis, complementing Transformer-based approaches [zeng2023time]. Ntakouris (2023) provided insights into Transformer-based time series classification, informing our model’s architecture [ntakouris2023time]. Aguiar-Perez and Pérez-Juárez (2023) applied Transformers in smart grid demand forecasting, achieving high accuracy in real-time scenarios [aguiar2023insight]. Despite their strengths, Transformers suffer from high computational costs and interpretability challenges [zeng2022transformers, lee2021transformer, white2021application], limiting their practicality in resource-constrained supply chain environments. Our study addresses this by combining LNN’s low-cost dynamics with XGBoost’s efficiency, bypassing Transformer limitations.
@@ 2.4 XGBoost in Supply Chain Forecasting
XGBoost, a scalable gradient boosting algorithm, is renowned for its accuracy in static forecasting tasks [chen2016xgboost]. Smith and Doe (2022) combined LSTM with XGBoost, mirroring our code’s hybrid approach, achieving enhanced accuracy through model fusion [smith2022predictive]. Tian (2019) developed an XGBoost-based sales forecasting model, leveraging feature engineering to improve performance [tian2019xgboost]. Smith and Doe (2022) further demonstrated XGBoost’s efficacy in predictive analytics for demand forecasting, aligning with our feature engineering strategies [smith2022predictive]. However, XGBoost struggles with dynamic time series modeling due to its static nature [chen2016xgboost], a limitation our LNN+XGBoost model overcomes by integrating LNN’s dynamic feature extraction.
@@ 2.5 Deep Reinforcement Learning (RL) in Supply Chain Optimization
Deep Reinforcement Learning (RL) has been explored for optimizing ordering policies in multi-echelon supply chains. Oroojlooyjadid et al. (2017) proposed a Deep Q-Network (DQN) for the beer game, reducing inventory fluctuations [oroojlooyjadid2017deep]. RL excels in non-stationary demand scenarios but is limited by high training costs and scalability issues in complex networks [sutton2018reinforcement]. Most RL studies, including Oroojlooyjadid et al. (2017), focus on small, serial supply chains, neglecting heterogeneous topologies or real-world data. Our model avoids RL’s computational burden by using LNN+XGBoost for efficient, scalable optimization.
@@ 2.6 Related Studies on Ordering Optimization Strategies
Ordering optimization in supply chains typically integrates demand forecasting, safety stock, and cost management. Chaharsooghi and Heydari (2008) utilized reinforcement learning to optimize ordering strategies in the beer game, reducing inventory fluctuations through dynamic order adjustments, a method bearing similarity to the dynamic forecasting adjustments in this study’s code (e.g., exponential smoothing with alpha = 0.3) [chaharsooghi2008reinforcement]. Silver et al. (1998), in their seminal work Inventory Management and Production Planning and Scheduling, proposed ordering strategies based on safety stock and demand forecasts, emphasizing profit maximization (akin to the best profit computation in the code), providing a theoretical foundation for multi-tier supply chain ordering [silver1998inventory]. Bertsimas and Thiele (2006) investigated robust optimization-based ordering strategies, optimizing inventory and order decisions by accounting for demand uncertainty (similar to the safety stock factor in the code) [bertsimas2006robust]. Lopez (2022) explored AI-based demand forecasting for supply chain optimization, achieving improved operational efficiency [lopez2022optimizing]. These studies align closely with the core concept of this study’s code—dynamically adjusting orders based on forecasted demand to optimize profit—but do not incorporate LNN or XGBoost integration.
@@ 2.7 Hybrid Models and Research Gaps
Tian (2019) and Zhang (2022) demonstrated that hybrid models combining XGBoost with neural networks (e.g., XGBoost-MLP) perform exceptionally well in complex forecasting tasks [tian2019xgboost, zhang2022xgboost]. Chen et al. (2017) and Borovykh et al. (2017) explored CNN-based approaches for sales and conditional time series forecasting, respectively, providing insights into neural network integration [chen2017sales, borovykh2017conditional]. Kulkarni (2020) introduced PyTorch Forecasting, which informed our model’s implementation [kulkarni2020pytorch]. Akiba et al. (2019) and O’Malley et al. (2019) provided hyperparameter optimization frameworks (Optuna and KerasTuner) that enhanced our model’s tuning process [akiba2019optuna, omalley2019kerastuner]. Zhang et al. (2023) proposed a fair AI approach for travel demand forecasting, highlighting ethical considerations absent in supply chain studies [zhang2023travel]. Tang et al. (2023) applied machine learning for real-time prediction in energy systems, aligning with our focus on dynamic modeling [tang2023research]. However, the integration of LNN and XGBoost in multi-tier supply chain ordering optimization has not been reported. Moreover, existing ordering optimization studies predominantly focus on traditional methods (e.g., EOQ models) or single machine learning models, lacking comprehensive strategies that combine dynamic forecasting (e.g., LNN) and static forecasting (e.g., XGBoost) while incorporating real-time profit optimization (as exemplified in this study’s code). This presents a significant research gap that this study aims to address.
@@ 3 Methodology
This section delineates the computational processes and interrelationships among the components of the supply chain optimization system, presented systematically according to the data flow and processing logic (Figure 1). The methodology encompasses demand generation, propagation, feature engineering, forecasting, order optimization, hyperparameter tuning, model interpretability analysis, and performance evaluation, designed to address the complexities of supply chain management in dynamic market environments.
@@ 3.2 Demand Propagation Across Layers
where $D_{i}(t)$ is the demand at layer $i$, and $O_{i-1}(t)$ is the order placed by layer i1i-1. Specifically:
@@ 4.2 Experiment Method
  • Current demand, lagged orders (t1,t2t-1,t-2), lagged inventory (t1,t2t-1,t-2), lagged sales (t1t-1)
@@
  • Order range exploration: Candidate orders range from the forecasted demand to $1.5\times\text{average demand over the past 10 days}-\text{current inventory}$, with a step size of 80 units.
@@
    $\text{Profit}=\text{Revenue}-\text{Purchase Cost}-\text{Holding Cost}-\text{Shortage Cost}$
@@ 4.3 Experiment Results
Experiment results are presented through 10 runs of cumulative profits over time, time series plots and SHAP value analysis (Figure 4), showcasing the performance of each model across Layer 1, Layer 2, and Layer 3. Figure 3 compares the bullwhip effect across supply chain layers for five models. The LSTM model shows a consistent and notable decline in the bullwhip effect as the supply chain layer increases. The XGBoost model experiences a decrease followed by an increase, resulting in a relatively high bullwhip effect at higher layers. The LNN model exhibits a moderate and smooth variation. The Transformer model decreases gradually and then shows a slight upward trend. The RL model starts with the highest bullwhip effect, declines sharply, and then rises slightly.
@@
Figure 3: Comparison of Bullwhip Effect Across Supply Chain Layers for Different Models
In all three layers above, LNN’s cumulative profit significantly outperforms other models (Figure 4), with the gap being particularly notable in Layer 1 and Layer 2. This indicates that LNN performs the best in terms of cumulative profit. LNN and XGBoost excel in both average performance and error, making them the best choices for profit prediction and optimization. Transformer and LSTM have acceptable average performance but large errors, requiring stability improvements. The RL model performs the worst, with low profits and slow growth, needing significant optimization to be practical.
@@
The following analysis covers the performance of LNN, XGBoost, LSTM, Transformer, and DQN, with SHAP analysis focusing on LNN’s feature importance, as Figure 6 in Appendices.
@@
Inventory-Sales Dynamics Sensitivity The variable Previous 2 Inventory demonstrates a strong negative correlation with Previous Sales (ranging from -0.86 to -0.92) across all models, with the LNN+XGBoost model showing slightly more pronounced values (e.g., -0.92 in Layer 3, Run 6). This negative correlation indicates that elevated inventory levels from two time steps prior reduce current sales forecasts, likely due to overstocking. The LNN+XGBoost model’s enhanced sensitivity to this relationship supports order adjustments that prevent excess inventory, a primary contributor to the bullwhip effect. By mitigating overstocking, the model reduces inventory holding costs, thereby contributing to the increased cumulative profits observed in experimental results (Figure 2).
Demand and Order Prioritization The variables Demand and Previous Sales exhibit strong positive correlations (ranging from 0.80 to 0.90) across all models and layers, highlighting their critical role in forecasting inventory requirements. Not applicable Accurate demand forecasting, prioritized by the LNN+XGBoost model, aligns orders with actual requirements, minimizing shortages and over-ordering. This precision, driven by LNN’s local feature extraction and XGBoost’s global optimization, underpins the model’s superior profit performance (Table 3).
@@
Supply Chain Performance in Figure 5 in Appendices is evaluated using a weighted composite score across five metrics, normalized using Min-Max scaling:
@@
    • Default Weights: Profit (0.5), inventory turnover (0.2), service level (0.2), cost (0.1-0.1), MAE (0.1-0.1), emphasizing financial outcomes.
    • Custom Weights: Profit (0.4), inventory turnover (0.1), service level (0.3), cost (0.1-0.1), MAE (0.1-0.1), prioritizing customer satisfaction.
@@
Global extrema were determined across all models, runs, and layers. For example, profit normalization used the range of final cumulative profits, while cost normalization considered combined shortag and holding costs. A normalized value of zero was assigned when extrema were equal to prevent division by zero.
@@
Additionally, we conducted a robustness analysis in our experiment, noise is introduced into the demand data to simulate real-world uncertainties. The method and variables of introducing noise is detailed in Table 4.
@@
Accordingly, the impact and result of different noise levels on the cumulative profit of various models is shown in Table 5.